OpenStudy (ria23):

x^2+y^2-8x+10y+15=0 I need to convert this to standard form, but I'm not quite sure how. . .

OpenStudy (jtvatsim):

OK, and what should standard form look like when we are done? What does "standard form" mean?

OpenStudy (ria23):

ax^2+bx+c I'm not sure how to answer what it means... It means like... it's in the standard form? .-.

OpenStudy (jtvatsim):

No worries you answered just fine. :) Actually the standard form you just gave is the standard form of a parabola (you use this when you have only x's). But we have both x's and y's. So we need to use a different standard form. I'll give it to you, but does that make sense so far?

OpenStudy (cwrw238):

you need to convert it to the standard form of a circle

OpenStudy (ria23):

Yea, it does, I knew there were different forms of it, but I wasn't sure which one...

OpenStudy (jtvatsim):

Right. So, since we have positive x^2 and positive y^2 in our equation, we will use try to make it look like the standard form of a circle. (x - a)^2 + (y - b)^2 = r^2

OpenStudy (anonymous):

I thought standard form was Ax + By = C ;o

OpenStudy (jtvatsim):

There are many "standard forms" depending on what shape your equation represents.

OpenStudy (jtvatsim):

Line: Ax + By = C Parabola: Ax^2 + Bx + C = 0 Circle: (x - A)^2 + (y - B)^2 = R^2 and so on...

OpenStudy (jtvatsim):

Of course, the question is "How do we get the mess we have into standard form?" :)

OpenStudy (jtvatsim):

Let's look at what we are given: x^2 + y^2 - 8x + 10y + 15 = 0

OpenStudy (jtvatsim):

Here's the first part of the trick. We move the +15 (just a number) to the right side of the equation like this... x^2 + y^2 - 8x + 10y = -15

OpenStudy (jtvatsim):

Is that OK so far?

OpenStudy (ria23):

Yep c: That's the one step I got and didn't know what to do next. n_n

OpenStudy (jtvatsim):

Cool. Now, we need to be a little sneaky... I am going to pair the x's and y's together so that it looks more organized like this... x^2 - 8x + y^2 + 10y = -15

OpenStudy (ria23):

Ok c:

OpenStudy (jtvatsim):

OK, now we will imagine how we can factor the x's and the y's... like this x^2 - 8x + y^2 + 10y = -15 (x )^2 + (y )^2 = -15 There's more just wait one sec. :)

OpenStudy (jtvatsim):

To get a -8x, I need to put a -4 in the factor like this x^2 - 8x + y^2 + 10y = -15 (x - 4 )^2 + (y )^2 = -15

OpenStudy (ria23):

If yhur trying to get a -8, and yhu put a -4 in and square it... won't yhu end up with a positive?

OpenStudy (jtvatsim):

Aha! you are right, we have to be careful... notice that (x - 4)^2 = (x - 4)(x - 4) = x^2 - 8x + 16 but we don't have a 16 to factor before! So, I need to add this 16 to the original equation... It sounds complicated but it just looks like this: x^2 - 8x + 16 + y^2 + 10y = -15 + 16 (x - 4)^2 + (y )^2 = -15 + 16

OpenStudy (jtvatsim):

I haven't done anything wrong. I just added 16 to both sides of the equation which is OK to do. Kind of scary, but does it make sense so far? We are almost done. :)

OpenStudy (ria23):

Yes it does. n_n

OpenStudy (jtvatsim):

OK, now the x's are all fixed. We only have to worry about the y's.

OpenStudy (jtvatsim):

But we can just use the same trick as before! :D

OpenStudy (jtvatsim):

I notice that if I put (y + 5)^2, I would get y^2 + 10y + 25...

OpenStudy (jtvatsim):

I really like the y^2 + 10y part, and I can fix the + 25 part easily, like this...

OpenStudy (jtvatsim):

x^2 - 8x + 16 + y^2 + 10y + 25 = -15 + 16 + 25 (x - 4)^2 + (y + 5)^2 = -15 + 16 + 25

OpenStudy (jtvatsim):

We got it! All that remains is to add the numbers on the right together to get the final answer.

OpenStudy (jtvatsim):

(x - 4)^2 + (y + 5)^2 = 26 The standard form is found! Whew... any questions?

OpenStudy (ria23):

:O No. Thank yhu for walking me through that. That was really helpful. n_n I have one more like this, do yhu mind if i try it out and post it in a new thread and tag yhu?

OpenStudy (jtvatsim):

Of course, let me know what you get. :)

OpenStudy (ria23):

Alright. n_n Thank yhu again

OpenStudy (jtvatsim):

No problem!